OpenStudy (anonymous):

Determine the tangent vector, unit normal,and bionormal for r(t)= (t^2 )+(2t+1)(t^3)

OpenStudy (amistre64):

is r(t) in component form? you know, with the ijk s?

OpenStudy (amistre64):

if so; then r' = tangent r'' = normal r'xr'' = binormal

OpenStudy (amistre64):

and any vector divided by its own magnitude goes unit

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

what if it ask me to calculate at the point on the curve where t=1

OpenStudy (amistre64):

then take the derivatives; input t=1, and voila!

OpenStudy (amistre64):

id try to type up a better response, but at the moment trying so on this site is pointless

OpenStudy (anonymous):

im still confuse with the binormal

OpenStudy (amistre64):

with any luck that fixed some problems

OpenStudy (anonymous):

its good

OpenStudy (amistre64):

ill assumes this is in vector component form, since thats about the only way itll make sense: r(t)= < t^2, (2t+1)(t^3) > r'(t) = <2t, 2t^3 + 3t^2(2t+1)> r''(t) = <2, 12t^2 + 6t(2t+1)>

OpenStudy (amistre64):

if t=1; and you want units; plug in t=1 to determine the vectors; and since a cross works best with 3 components; lets make z=0 r(1)= < 1, 3, 0 > r'(1) = <2, 11, 0> r''(1) = <2, 30, 0>

OpenStudy (anonymous):

i got some thing different.

OpenStudy (amistre64):

unitT = r'/|r'| = <2,11,0>/sqrt{125} unitN = r''/|r''| = <2,30,0>/sqrt{904} the Binormal is just the TxN, or r'xr''

OpenStudy (amistre64):

well, since im no good at mindreading; feel free to present what youve done :)

OpenStudy (anonymous):

its question number 4

OpenStudy (amistre64):

vectors are not single digits; unless you wanna consider the number line itself as the vector space

OpenStudy (amistre64):

make sure we got the right r(t)

OpenStudy (amistre64):

(2t+1)(t^3) = 6t^4 + t^3 derived = 24t^3 + 3t^2 which is not what youve got on your paper

OpenStudy (amistre64):

opps, typo

OpenStudy (amistre64):

2t^4 + t^3 8t^3 + 3t^2

OpenStudy (amistre64):

what you posted was this r(t)= (t^2)i+(2t+1)(t^3)j or is it spose to be: r(t)= (t^2)i +(2t+1)j +(t^3)k ???

OpenStudy (anonymous):

r(t)= (t^2)i +(2t+1)j +(t^3)k

OpenStudy (amistre64):

ahh, well that does change things then :)

OpenStudy (amistre64):

r' = 2t i + 2 j + 3t^2 k |r'| = sqrt(4t^2 + 4 + 9t^4) T(1) = <2,2,3>/5 r'' = 2 i + 0 j + 6t k |r''| = sqrt{4+0+36t^2} N(1) = <1,0,6>/sqrt(40) the wolfram is being retarded with this and cant verify if ive made a mistake ....

OpenStudy (anonymous):

ive been using wolfham all day

OpenStudy (amistre64):

unit T is spose to be: \(\large \frac{r'}{|r'|}\), which in this case is coming up hideous unit N is defined as; \(\large \frac{T'}{|T'|}\), which is every bit as much of a pain

OpenStudy (amistre64):

the agnitudes are not playing nice and making a monster of trying to go that route

OpenStudy (amistre64):

i also read up that tangent and normals are spose to be perp, so they dot to zero; but i cant seem to get this thing to dot up regradless

OpenStudy (amistre64):

http://tutorial.math.lamar.edu/Classes/CalcIII/TangentNormalVectors.aspx this might be useful ...